# SLayR: Scene Layout Generation with Rectified Flow

### v1

@@ 5 Experiments
5.1 Training and Evaluation Data
We use the ADE20K dataset [52] that contains approximately 27K images and manual semantic segmentations for indoor and outdoor scenes and a rich collection of object arrangements. The captions reflect the scene category with no additional detail (e.g., “beach”, “lecture room”). From captions and the semantic segmentations, we derive our ground-truth layouts. We use only the top $30$ largest bounding boxes per sample, as this is the default maximum number of bounding boxes supported by InstanceDiffusion  [45], padding samples with fewer bounding boxes.
5.2 Baseline Layout Generation Models
LayoutTransformer. We adapt LayoutTransformer  [11] to our task by training it on our dataset, and enabling it to take a global conditioning as input. To this end, we supply a bounding box that covers the entire image, with the attached global prompt as the first token.
GPT4o. LLMs have been used in related layout generation tasks successfully. Since they have a much larger size and are trained on much more different data, we do not consider them a direct baseline, but interesting to compare to. We selected GPT4o due to its effective performance across a wide range of tasks  [29], and since it was also trained on images. Querying the model with prompts from prior work  [21, 8] yielded empty layouts in $2/3$ of cases, as these templates assume details of the objects to be placed, e.g. “a park with a bench and a tree”. Therefore, we adapted the LLM-grounded Diffusion [21] template that provides seven in-context examples and instructs the model to decompose a generic scene prompt into subcomponents by utilizing chain of thought (COT) reasoning  [46]. We adopt the layout notation format and hyperparameters from  [21].
No Layout Structure. To understand the effect of layout conditioning on image generation, we also run our image generation pipeline with a layout containing no bounding boxes and just a global prompt.
Fixed Layout. To create an adversarial dataset with nonsensical layouts, we also create data with a fixed layout by distributing three bounding boxes evenly-spaced horizontally across the image and randomly sampling labels for each box.
5.3 Layout and Image Sampling
ADE20k is a relatively small dataset, and the training data from classes with few samples can be too little. To reduce the effect from these classes, we use the $15$ highest represented categories and add in five randomly selected categories. For each evaluated model, we generate $30$ layouts for all $20$ selected prompts, and an image conditioned on each layout and corresponding global prompt. This collection of images is also feasible to assess with our human evaluation.
5.4 Visual Results
We provide a qualitative overview of the generated layouts and the final images  Fig. 3. We label bounding boxes using the technique described in  Sec. 3.3, with all object labels present in ADE20k as our word bank.
LayoutTransformer struggles with arranging objects spatially, leading to images with unclear structure that do not depict the global prompt. GPT4o creates plausible layouts, but it struggles in making a variety of arrangements and representing the distribution of possible layouts. Our method produces the most plausible and diverse images across a range of global prompts, including both indoor and outdoor scenes.
@@ 5.4 Visual Results
Figure 3: Qualitative comparison (Best viewed up close). Layout objects that are depicted in the generated image are highlighted and labeled. No Layout produces scenes of little variation in viewpoint. GPT4o’s layouts are sensible, but also lack in variety. Layout Transformer produces layouts with implausible arrangements of objects, leading to images which do not depict the global prompt accurately. Our method creates the most plausible and variable layouts, leading to images that are diverse and look realistic.
Model FID ($\downarrow$) KID ($10^{-2}$)($\downarrow$) CMMD ($\downarrow$) IS ($\uparrow$) Inference Time (s) ($\downarrow$) Added Parameters ($\downarrow$)
LayoutTransformer 0.440.440.440.440.44 0.940.940.940.940.94 $1.34$ 8.348.348.348.348.34 24.6124.6124.6124.6124.61 23M
GPT4o $0.94$ $0.99$ $1.34$ $7.87$ $110.69$ $>$1T
Fixed Layout $1.16$ $2.31$ 1.321.321.321.321.32 $8.08$ N/A N/A
No Layout $5.06$ $1.11$DeclareFontSeriesDefault[rm]bfb1.001.001.$0$DeclareFontSeriesDefault[rm]bfb8.508.508.508.50 N/A N/A
@@
Table 1: Traditional Metric Comparison. Our method performs best on FID and KID while images generated from no layout, while images generated with a fixed layout perform second-best on CMMD, and third-best on IS, but both performed poorly on FID and KID. This suggests that FID and KID are more reliable metrics for the quality of the generated conditioning layout, and that CMMD and IS are not conclusive. Out of all the effective layout generation methods, ours is also the fastest and introduces the fewest parameters.
@@
5.5 Generated Image, Size, and Speed Metrics
To evaluate the generated images, we apply the established metrics FID [12], KID  [3], CMMD [15], and IS [37]. While FID, KID, and CMMD compare the distribution of generated images with the ground-truth ones, IS measures the variety of the generated images combined with how well the images can be classified as an object from a fixed set. Since the conditioned image generator may itself introduce details that alter the layout, and we want to limit these effects, we establish the ground-truth images by running InstanceDiffusion on the ground-truth layouts.
To compare the runtimes, we measure the average time it takes for each model to generate a batch of $30$ layout samples on an Nvidia A6000 GPU. Our model and LayoutTransformer generate these samples in parallel. We do not have direct access to GPT4o, and can only query it sequentially through an API. We also record the approximate number of model parameters that our baselines add to a complete text-to-image pipeline, factoring out the weights required for InstanceDiffusion.
We provide a quantitative overview in  Tab. 1. Our method performs state-of-the-art on FID and KID scores. The No Layout and Fixed Layout perform the poorest on these metrics, which indicates that these metrics are robust in effectively measuring structural similarity to the ground-truth images. We observe that the IS and CMMD metrics were optimal when there was no underlying structure at all, with our adversarial Fixed Layout performing second-best on CMMD, and third-best on IS. This indicates that CMMD and IS are ineffective in measuring the effect of the conditioning structure. Because CMMD is a metric which measures the similarity of images in the CLIP feature space, and is meant to align with a text encoding of a corresponding image caption, this metric arguably summarizes the images into global prompts and tends to be agnostic to the image’s spatial structure. In addition, using No Layout will effectively create an image with semantic matching of the global prompt and undermines CMMD’s effectiveness in assessing the underlying layout.
While IS is often used in assessing image generation techniques, it has been criticized for being ineffective at measuring image quality  [2, 5]. Images with multi-object structure might be at a disadvantage, as it measures how well the image fits into one of the single-object classification categories of the Inception-v3 model [44]. Furthermore, while IS is used for judging the variety of generated images, we show in  Sec. 5.7 that our method creates more varied images than the methods which are performing well on IS. These inconclusive findings led us to construct a user study. It is important to note that out of the effective layout generation methods we tested ours has the lowest number of parameters by over a factor of $5$, and is $37\%$ faster than the next fastest competitor.
5.6 Numerical Metrics Results
We show the numerical results from  Sec. 4.1 in  Fig. 5. LayoutTransformer performs comparable to our method in the object numeracy score ($O_{\text{Num}}$) and outperforms us in the position variance, meaning that it produces a large variety ($\sigma^{2}_{\text{Pos}}$) of the right types of objects. However, Figure 3 clearly shows that the generated layouts are not plausible.
GPT4o performs poorly on $O_{\text{Num}}$, although it was likely trained on much more data. When we examined the layouts, we observe that it tends to remain within a much stricter set of possible object labels than our method or LayoutTransformer. While it does well on $l_{\text{Pos}}^{(1)}$, it does poorer than our method on $l_{\text{Pos}}^{(2)}$, indicating that it has a reasonable understanding of where objects should be placed in the scene, but falls short in modeling the correct object interactions. Finally, our Fixed Layouts perform poorly as expected, which demonstrates the accuracy of the metrics.
[…]
[width=]sections/results/figures/baseline_bar_graph_combined
Figure 5: Quantitative Comparison. A comparison of our metrics described in Sec. 4.1 on our method and the layout-generating baselines introduced in Sec. 5.2. Our method is close or outperforms the baselines on $O_{\text{Num}}$, $l_{\text{Pos}}^{(1)}$, and $l_{\text{Pos}}^{(2)}$. Although we are lower than LayoutTransformer on $\sigma^{2}_{\text{Pos}}$,  Fig. 3 clearly shows that the generated variety from Layout Transformer is not plausible. This is further supported by our human evaluation in  Sec. 5.7.
5.7 Human Evaluation
We show the results from our human evaluation study in  Fig. 6. Aligned with the observations from Fig. 3, LayoutTransformer produces the highest variety, but the generated images are the least plausible. GPT4o delivers a similar plausibility as our method but is significantly more limited in variety, while No Layout is the most plausible but least diverse. The latter reveals the important insight that pure text-to-image generative denoising diffusion models do capture significant variety, and that pipelines with intermediate explicit representations can outperform them. We observe that our method really provides the highest trade-off in plausibility and variety, indicating that it can accurately generate a diverse distribution of images, in addition to being much smaller than its competitors.
Figure 6: Human Survey Results. Our method offers a superior trade-off between plausibility and variety, while even being much smaller. The error bars indicate the standard errors of the means and show that the results are statistically significant.
@@ 5.8 Additional Model Features
Finally, we present additional features which our model offers, which make it appealing for use in an end-to-end image generation pipeline. In  Fig. 4, we show examples of our model’s performance in different partial layout generation settings. This feature gives users even more fine-grained control over the image generation process. Furthermore, we demonstrate how our pipeline allows for editing of generated images in  Fig. 7 without introducing any additional model components. This is accomplished through modifying the intermediate scene layout, and rerunning InstanceDiffusion with the original seed and global prompt. We observe that the edits are not totally isolated to the altered objects and leave it for future work to develop improved conditional diffusion models.
@@ 6 Conclusion
We have introduced a text-to-image generation pipeline with an intermediate human-interpretable and controllable layout representation. While other methods produce images of scenes that are either very implausible or of little variety, our method clearly performs well on both of these axes and improves upon the state of the art, while being substantially smaller and faster. Our results show that that plain end-to-end pipelines are limited in the variety that they can capture and that an interpretable intermediate representation can capture this variety much better, while maintaining a high plausibility. In addition, we have introduced a suite of metrics for image layouts and generated images that allow reliable evaluation and can be repeated by the community. Both of these contributions together give rise to exploring image generation pipelines with intermediate explicit representations more in the future. In further work, we will scale up the approach to larger training datasets and add more detailed intermediate phases with segmentation masks.
@@ 7 Acknowledgements
The supplementary material is structured as follows. First, we present the full details of the human evaluation study performed to judge the generation quality in Sec.  A. Next, we introduce details about our partial conditioning procedure in Sec. B. We further provide detailed information about training data and hyperparameters in Sec. C, an ablation study with different of image generators in Sec. D, discussions about LayoutTransformer and GPT4o temperatures in Sec. E and F, respectively, and the GPT4o query template in Sec. G. Last, we provide a comparison between flow matching and DDIM in Sec. H and high-resolution results in Sec. I and Sec. J.
Appendix A Human Evaluation Details
Study Goal. Given the inconclusive results of the traditional metrics from Tab. 1, we aim to establish an effective metric for the quality of the generated layouts with respect to their effect on the final images. For the most accurate assessment, we provide a human-evaluation study that can be repeated by others.
In general, we expect that our generator should generate layouts that appear plausible and are also of a large variety. However, assessing human opinions for these criteria directly on layouts is challenging: the evaluators require time to understand the layout diagrams and explain them, and furthermore, assessments are hard to make without actually seeing the image. To make the study effective, we measure the effect of our model on the downstream generated images. Image qualities that are assessed in other studies (for example the overall quality and aesthetic appeal of the image in Liang et al. [22]) are highly dependent on the conditioned image generator. Therefore we consider these misleading for our case and introduce a suitable study in the following.
@@ Appendix A Human Evaluation Details
We assess the plausibility and variety of generated images that are conditioned on layouts from our method and three of the baselines presented in  Sec. 5.2: LayoutTransformer, GPT4o, and No Layout. We do not test against Fixed Layout, because this was constructed solely as an adversarial example for traditional metrics.
@@
Participant Selection and Ethics. Participants were recruited through Connect CloudResearch, a crowdsourcing service built on Amazon Mechanical Turk that implements rigorous quality control procedures to enhance the reliability of the participant pool in line with the study recommendations given by Otani et al. [30]. The study was approved by the Ethics Review Board of our institution, ensuring compliance with ethical standards. Prior to engaging in the tasks, all participants provided informed consent. The study was designed to be completed within $10$ minutes by each participant, who were compensated at an hourly rate of $13.02$ USD. This results in a total cost of 163 USD per run. Participants were anonymized, and we did not collect any personally-identifiable information.
@@
Appendix B Implementation of Partial Conditioning
We explain our adaptation of the RePaint [26] technique mentioned in  Sec. 3.3, which was used for the partial layout conditioning examples in  Fig. 4, An overview is presented in  Algorithm 1. The idea is that at every timestep, the intermediate sample $\mathbf{x}_{i}(t)$ is first updated with the rate of change provided by our model ($v$). Then the sample is slightly adjusted to conform to a path which will yield the values of the partial conditioning layout $\mathbf{y}_{i}$ at non-null entries after inference. The adjustment per iteration is small, allowing the model to obey the conditioning and plausibly completing the rest of the layout.
To fully understand the algorithm, we provide additional details by first defining notation: The partial layout representation $\mathbf{y}_{i}=\{\mathbf{y}_{i}^{j}\}_{j\in J}$ is defined like the layout representation in  Sec. 3.2 extended by null values $\varnothing$, a placeholder value for entries of $\mathbf{y}_{i}$ tokens where no conditioning is provided. To give an example, consider a partial conditioning layout where the token $\mathbf{y}_{i}^{j}$ enforces that a bounding box with the label chair must be present in the final layout, but can have any coordinates. We set ${\mathbf{b}_{i}^{j}}=(\varnothing,\varnothing,\varnothing,\varnothing)$, ${\mathbf{c}_{i}^{j}}$ to be the PCA-reduced CLIP embedding of the word chair and $\alpha_{i}^{j}=1$, and write:

### v2

@@ 5 Experiments
LayoutDiffusion Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 4: Qualitative comparison (Best viewed up close). Layout objects that are depicted in the generated image are highlighted and labeled. From a visual inspection, having no layout produces scenes of little variation in content. LayoutFlow’s layouts do not appear to capture scene structure. GPT4o’s layouts lack variety. Layout Transformer produces layouts with implausible arrangements of objects, leading to images which do not depict the global prompt accurately. Our method creates plausible and varied layouts, leading to images that are diverse and look realistic. These observations are supported by our human evaluation in  Fig. 5. Zoomed-in versions of these layouts for printing are available in the supplemental.
Model CMMD ($\downarrow$) FID ($\downarrow$) KID ($10^{-2}$) ($\downarrow$) VQA ($\uparrow$) HSPv2 ($\uparrow$) ImageReward ($\uparrow$)
LayoutDiffusion $0.61$ $1.08$ $1.92$ $0.34$ $0.19$ $-2.11$
LayoutFlow $0.25$ $0.80$ $0.88$ $0.80$ $0.23$ $-1.01$
LayoutTransformer $0.06$ $0.44$ $0.30$ $0.75$ $0.23$ $-1.00$
GPT4 $0.09$ $0.94$ $0.45$DeclareFontSeriesDefault[rm]bfb0.880.880.$8$DeclareFontSeriesDefault[rm]bfb0.250.250.250.25 $-0.51$
@@
Table 1: Image Metrics Comparison We evaluate traditional metrics and compare the images generated from layouts of different layout generators. To avoid biases of the image generator, we show the best score among the layout-to-image genators InstanceDiffusion  [46], GLIGEN [22], BoxDiff [50], and LMD+ [23] for each layout generator. Our method can achieve optimal performance across all measured metrics and achieves state of the art.
5.1 Experimental Setting
For a fair comparison, we train all models on the ADE20K dataset [59], which contains approximately 27K images and ground-truth layouts for indoor and outdoor scenes and a rich collection of object arrangements. The sample captions reflect the scene category with no additional constraints (e.g., beach, lecture room). We use the top $30$ largest bounding boxes per sample, as this is the default maximum number of bounding boxes supported by InstanceDiffusion  [46], padding samples with fewer bounding boxes. For evaluation, we use the $15$ highest represented categories and add in five randomly selected categories to include the dataset’s long tail distribution. For each evaluated model, we generate $30$ layouts for all $20$ selected prompts, and an image conditioned on each layout and corresponding global prompt. The size of this collection of images is feasible to assess with our human evaluation.
We adapt LayoutTransformer  [10], LayoutDM  [14] and LayoutFlow  [9] from the UI generation setting by treating the global caption as a scene-wide bounding box and conditioning the model on this bounding box during inference. As we demonstrate in  Fig. 2, existing LLM-based approaches [23, 7] do not succeed off-the-shelf in our task. However, by adjusting the template of LLM-Grounded Diffusion [23] with relevant in-context-learning examples from ADE20K, we avoid degenerate solutions to have a fair comparison. For the underlying LLM, we select the recent large-scale LLM GPT4o [32].
5.2 Generated Image Evaluation
Text-to-Image Metrics. To evaluate the generated images, we apply the established image generation metrics CMMD [15], FID [11], KID [3], VQA  [26], HPSv2  [48], and ImageReward  [52]. CMMD, FID and KID compare the distribution of generated images with a ground-truth distribution, while VQA, HSPv2 and ImageReward assess general image quality and alignment with a global caption. Since the conditioned image generator may itself lead to biases in image generation quality, for CMMD, FID and KID we establish the ground-truth images by running the layout-to-image generator on the ground-truth layouts. Furthermore, as different image generators are also biased towards different layout generators, for each layout generator we choose the optimal score over the possible combinations of layout and image generator. The results are shown in Tab. 1. As visible, our method outperforms all other methods and achieves state of the art.
Human Evaluation. As shown in  Fig. 5, our model achieves a state-of-the-art balance in image plausibility and variety as well. We display the approximate number of model parameters added to the full text-to-layout-to-image pipeline by the layout generators that can be locally run. Our model is the smallest by over a factor $3$.
Visual Results. We provide a qualitative overview of the generated layouts and the final images in Fig. 4. Due to space limitations, we display the baselines that performed best in  Tab. 1. We label bounding boxes by querying with all text labels present within ADE20K. From a visual inspection, LayoutTransformer struggles with arranging objects spatially, leading to images with unclear structure that do not depict the global prompt. GPT4o layouts appear somewhat flat, while struggling to make a variety of layouts. Our method appears to produce both plausible and diverse images across a range of global prompts of indoor and outdoor settings.
Figure 5: Human Survey Results. Our method offers a superior trade-off between plausibility and variety across all measured layout-to-image generators, while being a much smaller model. The error bars indicate the standard errors of the means and show that the results are statistically significant.
@@ 5.2 Generated Image Evaluation
5.3 Layout Evaluation and Inference Speed
We evaluate our method on the layout metrics proposed in section Sec. 4.2 in Tab. 2. While LayoutTransformer outperforms our model on the variety measure $\sigma^{2}_{\text{Pos}}$, we observe in Fig. 4 that the layouts lack plausibility. For example, the floor in the leftmost example appears at the top and the ashcan on the rightmost example is significantly too large.
We also measure the average time it takes for each model to generate a batch of $30$ layout samples on an Nvidia A6000 GPU in Tab. 2. As we do not have direct access to GPT4o, we query it sequentially through an API. Our method ranks second in speed. Although LayoutFlow’s default settings enable faster inference, we observe no definitive improvement in its layout statistics when the number of inference steps are raised to match our model’s speed.
Model $O_{\text{Num}}$ $\sigma^{2}_{\text{Pos}}$ $l_{\text{Pos}}^{(1)}$ $l_{\text{Pos}}^{(2)}$ mIoU Time (s)
GPT4 3.71 93 4.17 1.42 0.10 111.0
LayoutDiffusion 2.12 65 1.47 0.71 0.00 138.0
LayoutFlow 3.01 142 1.48 0.72 0.01 0.5
LayoutFlow (More steps) 2.96 143 1.44 0.65 0.01 15.5
LayoutTransformer 0.90 231 3.09 1.21 0.15 25.0
Ours 1.14 187 4.76 1.93 0.17 15.5
Table 2: Layout Metrics, and Inference Speed. A comparison of our metrics introduced in Sec. 4.2. $O_{\text{Num}}$ is scaled by $10^{2}$. Our method achieves the best on mIoU, $l_{\text{Pos}}^{(1)}$, and $l_{\text{Pos}}^{(2)}$, and is behind LayoutTransformer in $O_{\text{Num}}$ and $\sigma^{2}_{\text{Pos}}$. The closer inspection in Sec. 5.2 reveals that LayoutTransformer falls short in terms of plausibility and image quality, indicating that it generates a large variety but implausible layouts.
5.4 Scaling to MSCOCO
MSCOCO contains approximately 125K samples that consist of an image, the annotated scene layout, and five alternative global captions that are full sentence descriptions (e.g. ”A black cat laying on top of a bed next to pillows”). We explore how our model scales to MSCOCO [25] by training on this dataset using the same architecture and hyperparameters as our main model. We provide some qualitative examples in Fig. 7 and show that we are able to generate appealing images for the complex prompts with our method.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: MSCOCO Trained Model: Layouts and generated images from the prompts: An image with two clocks, The food is ready to be eaten on the table, A group of teens playing frisbee in a park, and A car to the left of a bus
@@ 5.5 Additional Model Features
At last, we present the additional features that our model offers that make it appealing for use in an end-to-end image generation pipeline. In  Fig. 6, we show examples of our model’s performance in different partial layout generation settings. This feature gives users even more fine-grained control over the image generation process.
Furthermore, we demonstrate how our pipeline allows for editing of generated images in  Fig. 8. This is accomplished through modifying the intermediate scene layout, and rerunning InstanceDiffusion with the original seed and global prompt. We observe that the edits are not totally isolated to the altered objects and leave it for future work to develop improved conditional diffusion models.
@@ 6 Conclusion
We have introduced a text-to-image generation pipeline with an intermediate human-interpretable and controllable layout representation. With a substantially smaller model, we achieve state-of-the-art performance in generated image quality metrics among competing baselines, as well as state of the art when considering both of the axes plausibility and variety. In addition, we have introduced a suite of metrics for scene layouts and generated images for reliable evaluation. Both of these contributions together enable further exploration of image generation pipelines with intermediate explicit representations in the future.
@@ 7 Acknowledgements
The supplementary material is structured as follows. First, we present the full details of the human evaluation study performed to judge the generation quality in Sec.  8. Next, we introduce details about our partial conditioning procedure in Sec. 9. We further provide detailed information about training data and hyperparameters in Sec. 10, discussions about LayoutTransformer and GPT4o temperatures in Sec. 11 and 12, respectively, and the GPT4o query template in Sec. 13. Last, we provide a comparison between rectified flow and DDIM in Sec. 14 and high-resolution results in Sec. 15 and Sec. 16.
8 Human Evaluation Details
Study Goal. Although our method achieves optimal performance in  Tab. 1, we aim to confirm that these metrics, which were designed for measuring the quality of text-to-image models, are applicable to text-to-layout-to-image models. We also want to control for the effect which the layout-to-image model could have on the final quality, and assess how effective the underlying layouts are in the image generation process. To this end, we provide a human-evaluation study that can be repeated by others.
In general, we want a text-to-layout model to generate layouts that appear plausible and are also of a large variety. However, assessing human opinions for these criteria directly on layouts is challenging: the evaluators require time to understand the layout diagrams and explain them, and furthermore, assessments are hard to make without actually seeing the image. To make the study effective, we measure the effect of our model on the downstream generated images. Image qualities that are assessed in other studies (for example, the overall quality and aesthetic appeal of the image in Liang et al. [24]) are highly dependent on the conditioned image generator. Therefore, we consider these misleading for our case and introduce a suitable study in the following.
@@ 8 Human Evaluation Details
Participant Selection and Ethics. Participants were recruited through Connect CloudResearch, a crowdsourcing service built on Amazon Mechanical Turk that implements rigorous quality control procedures to enhance the reliability of the participant pool in line with the study recommendations given by Otani et al. [33]. The study was approved by the Ethics Review Board of our institution, ensuring compliance with ethical standards. Prior to engaging in the tasks, all participants provided informed consent. The study was designed to be completed within $15$ minutes by each participant, who were compensated at an hourly rate of $13.02$ USD. This results in a total cost of 245 USD per run. Participants were anonymized, and we did not collect any personally-identifiable information.
@@
9 Implementation of Partial Conditioning
We explain our adaptation of the RePaint [29] technique mentioned in  Sec. 3.3, which was used for the partial layout conditioning examples in  Fig. 6, An overview is presented in  Algorithm 1. At every timestep, the intermediate sample $\mathbf{x}_{i}(t)$ is first updated with the rate of change provided by our model ($v$). Then the sample is slightly adjusted to conform to a path which will yield the values of the partial conditioning layout $\mathbf{y}_{i}$ at non-null entries after inference.
Some additional algorithm notation: The partial layout representation $\mathbf{y}_{i}=\{\mathbf{y}_{i}^{j}\}_{j\in J}$ is defined like the layout representation in  Sec. 3.2 extended by null values $\varnothing$, a placeholder value for entries of $\mathbf{y}_{i}$ tokens where no conditioning is provided. To give an example, consider a partial conditioning layout where the token $\mathbf{y}_{i}^{j}$ enforces that a bounding box with the label chair must be present in the final layout, but can have any coordinates. We set ${\mathbf{b}_{i}^{j}}=(\varnothing,\varnothing,\varnothing,\varnothing)$, ${\mathbf{c}_{i}^{j}}$ to be the PCA-reduced CLIP embedding of the word chair and $\alpha_{i}^{j}=1$, and write: